OpenStudy (anonymous):

∫_(-1)^1 {dx/ (t ^(2)-25)} you dont have to solve i just want to know whats my u

OpenStudy (amistre64):

hyperbolic ?

OpenStudy (anonymous):

I suppose technically there is no "u" because it is slightly more complicated than a u-sub. You would have to do trig sub if I'm seeing it right.

OpenStudy (amistre64):

can you z sub it? like you gotta use a z lol

OpenStudy (anonymous):

\[\int\limits_{0}^{4}{\frac{2t}{t^2-25}}dt\] sorry this is the corect question

OpenStudy (amistre64):

t = z^(1/2) or so?

OpenStudy (anonymous):

u sub

OpenStudy (amistre64):

thats just an ln

OpenStudy (amistre64):

the top is the derivative of the bottom; it been ln-ized

OpenStudy (anonymous):

ok...i dont understand what your saying... amistre

OpenStudy (amistre64):

\[\int\frac{dx}{x}\implies ln|x|+c\]

OpenStudy (amistre64):

in this case; x = t^2-25; and dx = 2t

OpenStudy (anonymous):

1/x yes....but i dont see it in that integral

OpenStudy (anonymous):

Yeah, you can use u=t^2-25. Then du=2t Then your integral becomes \[\int\limits \frac{du}{u}\] You need to re-evaluate your limits of integration however.

OpenStudy (amistre64):

whenever the top is the derivative of the bottom; you got i log intergral

OpenStudy (anonymous):

we didnt do limits of integration

OpenStudy (amistre64):

no need to go thru the extra work; ln(t^2-25) is your integral

OpenStudy (amistre64):

[ln(x^2)]' = 2x/x^2 ; the top is the derivative of the bottom thanks to the chain rule

OpenStudy (anonymous):

You have the integral from 0 to 4. Those are your limits.

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

ln|t^2-25| then ;)

OpenStudy (anonymous):

okay thank you both

OpenStudy (anonymous):

ok i dont see ow to get the answe

OpenStudy (amistre64):

the answer is just in recognizing the integral as an adaptation to \(\int\frac{dx}{x}\)

OpenStudy (anonymous):

then?

OpenStudy (amistre64):

then you subtract the ln|t^2 -25| using the upper and lower bounds F(upper) - F(lower)

OpenStudy (amistre64):

\[ln|(4^2)-25|-ln((0)^2-25)=?\]

OpenStudy (amistre64):

there spose to be absolute value signs around the ln parts

OpenStudy (amistre64):

ln(9) - ln(25) = ?

OpenStudy (amistre64):

ln(9/25) = ?

OpenStudy (anonymous):

i thought u cant find the ln of a negative number

OpenStudy (amistre64):

you cant; which is why you absolute value it to make it positive

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

the books show the correct way to do it is with: \[ln|x|\] and then give a reason why its good either way, but i cant recall the details :)

OpenStudy (anonymous):

okay thank you wow i understand now!!!

OpenStudy (amistre64):

:) youre welcome